# Chirality, confinement and dimensionality govern re-entrant transitions in active matter

### v1

@@ Abstract
The non-equilibrium dynamics of individual chiral active particles underpin the complex behavior of chiral active matter. Here we present an exact analytical framework, supported by simulations, to characterize the steady states of two-dimensional chiral active Brownian particles and three-dimensional torque-driven counterparts in a harmonic trap. Using a Laplace-transform approach of the Fokker–Planck equation, we derive closed-form expressions for displacement moments and excess kurtosis, providing a precise probe of non-Gaussian statistics. Our analysis reveals three distinct regimes: bimodal active states with off-center peaks, Gaussian-like passive states, and weakly heavy-tailed distributions unique to two dimensions. We show that dimensionality plays a decisive role: in two dimensions, increasing chirality suppresses activity and restores passive behavior, while in three dimensions torque preserves activity along the torque axis, producing anisotropic steady states. These behaviors are captured by simple active length-scale arguments that map the boundaries between passive and active phases. Our results offer concrete experimental signatures — including kurtosis crossovers, off-center peaks, and torque-induced anisotropy — that establish confinement as a powerful tool to probe and control chiral and torque-driven active matter.
@@ I Introduction
Active matter remains out of equilibrium by converting energy into self-propulsion, producing persistent motion and breaking time-reversal symmetry [1, 2, 3, 4, 5, 6, 7]. Natural active systems include motor proteins, bacteria, and cells, as well as larger collectives such as fish schools and bird flocks [8, 9, 10, 11, 12]. Artificial realizations range from self-propelled colloids and synthetic microswimmers to vibrated polar disks and robotic swarms [13, 14, 15, 16, 17]. With intrinsic chirality or an applied torque, active matter forms a distinct class of non-equilibrium systems where self-propulsion, coupled with angular velocity, gives rise to a wide range of complex behaviors [18, 19, 20, 21, 22, 23, 24, 25, 26].
When chirality is present, an active particle traces circular or helical paths, set by an intrinsic angular velocity that often comes from asymmetries in its shape or propulsion mechanism [27, 28, 29, 30, 31, 32]. In odd active matter, chirality can arise spontaneously, giving rise to striking properties such as odd viscosity, odd elasticity, and odd diffusivity [33, 34, 35, 36]. Chiral motion has been observed in several experimental contexts, such as in active biomolecules [37], motor proteins [38], microtubules [39], cells and bacteria [40, 41, 42, 43, 44, 45, 46, 47]. In particular, bacteria swimming near a substrate, propelled by flagellar rotation, display clear chiral trajectories [46, 47]. On the theoretical side, chiral motion has been analyzed through mean trajectories in both two- and three-dimensional systems [48, 49, 50, 27, 51, 52, 53, 54, 55, 46, 47, 56, 57].
@@
Using an exact analytical framework and numerical simulations, we present a comprehensive investigation of the non-equilibrium steady-state (NESS) properties of two-dimensional chiral active Brownian particles (ABPs) and three-dimensional torque-driven ABPs confined within a harmonic potential. We employ a Laplace transform-based method, originally developed for semiflexible polymers [63, 64] and more recently extended to active particle systems [65, 66, 67, 68, 69], to use the Fokker-Planck equation to derive exact analytical expressions for arbitrary moments of key observables, including the second and fourth moments of displacement and the excess kurtosis. In particular, the excess kurtosis emerges as a sensitive probe of non-equilibrium behavior in the steady state, capturing how activity, chirality or torque, and trap strength together drive deviations from equilibrium. Building on prior observations in non-chiral active Brownian particles (ABPs), we show that the inclusion of chiral motion significantly modifies the nature of the re-entrant transition between passive and active states under increasing trap strength. [70, 66].
Figure 1: Schematic illustration of torque in two and three dimensional active particles, showing their typical steady state trajectories under harmonic confinement.
@@ II A trapped Chiral ABP in 2D
Here $\mu=k/\gamma$ denotes the scaled trap strength with respect to the viscous drag coefficient $\gamma$. The translational Wiener process $d\bm{B}$ and the rotational process $dW$ both have zero mean, with variances $\langle dB_{i}dB_{j}\rangle=\delta_{ij},dt$ (for $i,j\in{x,y}$) and $\langle dW^{2}\rangle=dt$. The corresponding diffusion coefficients $D$ (translational) and $D_{r}$ (rotational) lead to the units of time and length in terms of the persistence time $\tau_{r}=D_{r}^{-1}$ and length scale $\ell=\sqrt{D/D_{r}}$. The dimensionless control parameters are the Péclet number ${\rm{Pe}}=v_{0}/\sqrt{DD_{r}}$ (activity), chirality $\Omega=\omega/D_{r}$, and trap strength $\beta=\mu/D_{r}$. The dimensionless position and time will be denoted by $\tilde{{\textbf{r}}}={\textbf{r}}/\ell$ and $\tilde{t}=t/\tau_{r}$.
The harmonic trap guarantees an asymptotic steady state, which we analyze using both exact analytic methods and numerical simulations. We numerically integrate the dimensionless version of the Langevin equations using the Euler-Maruyama scheme, comparing the results with our analytic predictions. The steady-state moments are derived by formulating the Fokker-Planck equation, applying a Laplace transform, and utilizing the resulting moment-generating equations. Key findings are summarized below, and the full analytic derivations are provided in the Supplementary Material [71].
@@ II.1 Lower order moments at steady state
The presence of the harmonic trap ensures a steady state with $\langle{\tilde{\textbf{r}}}{\rangle}_{\rm st}=0$. In equilibrium, the position distribution $p({\tilde{\textbf{r}}})=\exp{\left(-{{\tilde{\textbf{r}}}^{2}}/{2\sigma^{2}}\right)}/{\sqrt{2\pi\sigma^{2}}}$ is Gaussian, where $\sigma^{2}=\langle{\tilde{\textbf{r}}}^{2}{\rangle}_{\rm st}/d$ in $d$-dimensions with the mean-squared displacement (MSD) given by $\langle{\tilde{\textbf{r}}}^{2}{\rangle}_{\rm st}=d/\beta$. This form is expected to hold at low enough activity. However, at large activity, even for non-chiral ABPs, the distribution deviates significantly from Gaussian behavior. These non-Gaussian departures can be characterized by exact calculations of fourth moment and kurtosis of displacement [66]. We compare the numerically obtained probability distributions with the Gaussian form supplemented by the exact MSD estimate shown below, to highlight the non-Gaussian departures (see solid lines in Figs. 2(d,e,f)).
In 2D, the distributions of non-chiral ABPs exhibited a circularly symmetric ring around the trap minimum, with particles climbing the trap potential until reaching a radial extent determined by the force balance condition: $\tilde{\rm{r}}^{0}_{\rm ac}={\rm Pe}/\beta$ [70, 66]. An equivalent estimate for chiral ABPs can be derived by analyzing the deterministic part of Langevin equations: $\dot{\tilde{x}}={\rm Pe}\cos\phi-\beta\tilde{x}$, $\dot{\tilde{y}}={\rm Pe}\sin\phi-\beta\tilde{y}$ and $\dot{\phi}=\Omega$. Solving these equations and obtaining the time-asymptotic expressions for $\tilde{x}(t)$ and $\tilde{y}(t)$, we find $\tilde{\rm{r}}_{\rm ac}={\rm Pe}/\sqrt{\beta^{2}+\Omega^{2}}$ (see Supplementary Material [71]). As we show later, this estimate agrees well with the off-center maxima of the probability distributions at high activity.
By applying the Fokker-Planck equation method described in the previous section, the first result we obtain is the MSD at steady state,
@@
The first term, denotes the equilibrium contribution in 2D, and the second term is entirely due to activity which increases as ${\rm Pe}^{2}$, but decreases quadratically with chirality as $\Omega^{-2}$.
@@
II.2 Excess kurtosis and phase diagram
In the steady state, the deviation from Gaussian behavior is quantified by the excess kurtosis, defined as $\tilde{\cal{K}}_{\rm st}={\langle{{\tilde{\textbf{r}}}^{4}}\rangle_{\rm st}}/{\tilde{\mu}_{4}^{\rm st}}-1$, with $\tilde{\mu}_{4}=\langle{\tilde{\textbf{r}}}^{2}\rangle^{2}+2\langle\tilde{r}_{i}\tilde{r}_{j}\rangle^{2}$ giving the estimate of the fourth moment in a Gaussian process. In 2D, $\tilde{\mu}_{4}^{\rm st}=2\langle{\tilde{\textbf{r}}}^{2}\rangle_{\rm st}^{2}$ gives the exact expression (see Supplementary Material [71])
@@ II.2 Excess kurtosis and phase diagram
with $\tilde{\cal{K}}_{\rm st}$ vanishing linearly with $\beta$. In the small $\beta$ limit, $\tilde{\cal{K}}{\rm st}$ behaves differently depending on $\Omega$: it is positive when $\Omega^{2}>7/5$ and negative when $\Omega^{2}<7/5$. As illustrated in Fig. 2(a), for instance, at $\Omega=5$ (blue square symbols) where $\Omega^{2}>7/5$, $\tilde{\cal{K}}{\rm st}$ is initially positive at small $\beta$, but then becomes negative, following a $\beta^{2}$ dependence as the trap stiffness increases. After reaching a minimum at intermediate values of $\beta$, it begins to rise again, ultimately vanishing as $\sim-\beta^{-2}$ in the $\beta\to\infty$ limit.
@@
Phase diagram: Fig. 2(c) shows a phase diagram in the $\Omega-\beta$ plane, highlighting the transition between passive (equilibrium-like) and active phases based on deviations from Gaussian statistics. While there is no true phase transition in this single-particle system, the diagram reveals a re-entrant crossover, shifting from passive to active and back to passive phases as trap stiffness increases. We identify three phases based on steady-state excess kurtosis: negative kurtosis defines the active phase (red) with a symmetric bimodal distribution with off-center peaks; vanishing kurtosis corresponds to the passive phase (blue) with a Gaussian-like distribution; and small positive kurtosis indicates another active phase (green) with a weakly heavy-tailed distribution, unique to chiral ABPs and absent in trapped achiral ABPs [66]. The solid line represents $\tilde{\cal{K}}_{\rm st}=0$, separating negative and positive kurtosis.
@@
Probability distributions: In Figs.2(d-f), we present radial position distributions for the parameter values marked by Roman and Indo-Arabic numerals in the phase diagram of Fig.2(c). The solid lines correspond to the Gaussian distribution $\rm p({\tilde{\textbf{r}}})=\exp{\left(-{{\tilde{\textbf{r}}}^{2}}/{\langle{\tilde{\textbf{r}}}^{2}\rangle_{\rm st}}\right)}/{\sqrt{\pi\langle{\tilde{\textbf{r}}}^{2}\rangle_{\rm st}}}$ using $\langle{\tilde{\textbf{r}}}^{2}\rangle_{\rm st}$ for the corresponding parameter values. These help identify deviations from passive, equilibrium-like behavior.
@@
In Fig. 2(d), the position distribution is shown for varying trap stiffness $\beta=0.08~\mathrm{(i)},0.33~\mathrm{(ii)},10~\mathrm{(iii)},2\times 10^{4}~\mathrm{(iv)}$ with constant $\Omega=5$. The transition is observed from a weakly heavy-tailed active state (${\cal K}_{\rm st}>0$) to a Gaussian passive state, then to an active state with off-center peak (${\cal K}_{\rm st}<0$), and finally back to a Gaussian passive state. Notably, for $\mathrm{(i)}~\beta=0.08$, although the excess kurtosis is positive, it is small ($\tilde{\cal{K}}_{\rm st}\approx 0.0084$), and the distribution appears nearly Gaussian. For $\mathrm{(ii)}~\beta=0.33$ on the zero excess kurtosis line, a Gaussian distribution is expected and observed. At $\mathrm{(iii)}~\beta=10$, a peak in the distribution away from the trap center, at $\tilde{\rm{r}}_{\rm ac}={\rm Pe}/\sqrt{\beta^{2}+\Omega^{2}}$, indicates an active state. At $\mathrm{(iv)}~\beta=2\times 10^{4}$, the distribution reverts to Gaussian.
@@
In Fig. 2(e), radial distributions across different chirality values ($\Omega$) with $\beta=10$ are shown, revealing transitions from active (off-center peak at $\tilde{\rm{r}}_{\rm ac}$) to passive (Gaussian) states. Fig. 2(f) at $\beta=0.15$ shows radial distributions at various $\Omega$ corresponding to small positive or negative kurtosis values with small departures from Gaussian.
@@
Analysis of the phase diagram: The phase diagram reveals an intermediate regime of $\beta$ supporting active dynamics marked by negative kurtosis. For $\beta>1$, persistence dominates over trap relaxation time allowing for active phase. Neglecting diffusion, the maximal radial reach of a trapped particle is $\tilde{\rm{r}}_{\rm ac}={\rm{Pe}}/\sqrt{\beta^{2}+\Omega^{2}}$, while translational noise broadens the distribution over $\ell_{\rm eq}=\sqrt{2/\beta}$. The off-center peak becomes negligible when $\tilde{\rm{r}}_{\rm ac}\lesssim\ell_{\rm eq}$, solution of the quadratic equation at $\tilde{\rm{r}}_{\rm ac}=\ell_{\rm eq}$ defines two bounds of the active phase,

### v2

@@ Abstract
The non-equilibrium dynamics of individual chiral active particles underpin the complex behavior of chiral active matter. Here we present an exact analytical framework, supported by simulations, to characterize the steady states of two-dimensional chiral active Brownian particles and three-dimensional torque-driven counterparts in a harmonic trap. Using a Laplace-transform approach to the Fokker–Planck equation, we derive closed-form expressions for displacement moments and excess kurtosis, providing a precise probe of non-Gaussian statistics. Our analysis reveals three distinct regimes characterized by bimodal distributions with off-center peaks, Gaussian-like distributions, and weakly heavy-tailed distributions unique to two dimensions. We show that dimensionality plays a decisive role: in two dimensions, increasing chirality suppresses activity and restores Gaussian-like behavior, whereas in three dimensions, torque sustains anisotropic steady states and preserves non-Gaussianity even at high chirality. These behaviors are captured by simple active length-scale arguments that map the boundaries between Gaussian-like and non-Gaussian states. Our results offer concrete experimental signatures — including kurtosis crossovers, off-center peaks, and torque-induced anisotropy — that establish confinement as a powerful tool to probe and control chiral and torque-driven active matter.
@@ I Introduction
Active matter remains out of equilibrium by converting energy into self-propulsion, producing persistent motion and breaking time-reversal symmetry [1, 2, 3, 4, 5, 6, 7]. Natural active systems include motor proteins, bacteria, and cells, as well as larger collectives such as fish schools and bird flocks [8, 9, 10, 11, 12]. Artificial realizations range from self-propelled colloids and synthetic microswimmers to vibrated polar disks and robotic swarms [13, 14, 15, 16, 17]. With intrinsic chirality or an applied torque, active matter forms a distinct class of non-equilibrium systems where self-propulsion, coupled with angular velocity, gives rise to a wide range of complex behaviors [18, 19, 20, 21, 22, 23, 24, 25, 26, 27].
When chirality is present, an active particle traces circular or helical paths, set by an intrinsic angular velocity that often comes from asymmetries in its shape or propulsion mechanism [28, 29, 30, 31, 32, 33]. In odd active matter, chirality can arise spontaneously, giving rise to striking properties such as odd viscosity, odd elasticity, and odd diffusivity [34, 35, 36, 37]. Chiral motion has been observed in several experimental contexts, such as in active biomolecules [38], motor proteins [39], microtubules [40], cells and bacteria [41, 42, 43, 44, 45, 46, 47, 48]. In particular, bacteria swimming near a substrate, propelled by flagellar rotation, display clear chiral trajectories [47, 48]. On the theoretical side, chiral motion has been analyzed in both two- and three-dimensional systems [49, 50, 51, 28, 52, 18, 53, 54, 55, 47, 48, 56, 57, 58].
@@
Using an exact analytical framework and numerical simulations, we present a comprehensive investigation of the non-equilibrium steady-state (NESS) properties of two-dimensional chiral active Brownian particles (ABPs) and three-dimensional torque-driven ABPs confined within a harmonic potential. We employ a Laplace transform-based method, originally developed for semiflexible polymers [64, 65] and more recently extended to active particle systems [66, 67, 68, 69, 70], to use the Fokker-Planck equation to derive exact analytical expressions for arbitrary moments of key observables, including the second and fourth moments of displacement and the excess kurtosis. In particular, the excess kurtosis emerges as a sensitive probe of non-equilibrium behavior in the steady state, capturing how activity, chirality or torque, and trap strength together drive deviations from equilibrium. Building on prior observations in non-chiral active Brownian particles (ABPs), we show that the inclusion of chiral motion significantly modifies the nature of the re-entrant transition between Gaussian and non-Gaussian states under increasing trap strength. [71, 67].
Figure 1: Schematic representation of torque effects in two- and three-dimensional active particles, illustrating their typical steady-state trajectories under harmonic confinement.
@@ II A trapped Chiral ABP in 2D
Here $\mu=k/\gamma$ denotes the scaled trap strength with respect to the viscous drag coefficient $\gamma$. The translational Wiener process $d\bm{B}$ and the rotational process $dW$ both have zero mean, with variances $\langle dB_{i}dB_{j}\rangle=\delta_{ij}dt$ (for $i,j\in{x,y}$) and $\langle dW^{2}\rangle=dt$. The corresponding diffusion coefficients $D$ (translational) and $D_{r}$ (rotational) lead to the units of time and length in terms of the persistence time $\tau_{r}=D_{r}^{-1}$ and length scale $\ell=\sqrt{D/D_{r}}$. The dimensionless control parameters are the Péclet number ${\rm{Pe}}=v_{0}/\sqrt{DD_{r}}$ (activity), chirality $\Omega=\omega/D_{r}$, and trap strength $\beta=\mu/D_{r}$. The dimensionless position and time will be denoted by $\tilde{{\textbf{r}}}={\textbf{r}}/\ell$ and $\tilde{t}=t/\tau_{r}$.
The harmonic trap guarantees an asymptotic steady state, which we analyze using both exact analytic methods and numerical simulations. We numerically integrate the dimensionless version of the Langevin equations using the Euler-Maruyama scheme, comparing the results with our analytic predictions. The steady-state moments are derived by formulating the Fokker-Planck equation, applying a Laplace transform, and utilizing the resulting moment-generating equations. Key findings are summarized below, and the full analytic derivations are provided in the Appendix-A.
@@ II.1 Lower order moments at steady state
The presence of the harmonic trap ensures a steady state with $\langle{\tilde{\textbf{r}}}{\rangle}_{\rm st}=0$. In equilibrium, the position distribution $\textrm{p}({\tilde{\textbf{r}}})=\exp{\left(-{{\tilde{\textbf{r}}}^{2}}/{2\sigma^{2}}\right)}/{\left(2\pi\sigma^{2}\right)^{d/2}}$ is Gaussian, where $\sigma^{2}=\langle{\tilde{\textbf{r}}}^{2}{\rangle}_{\rm st}/d$ in $d$-dimensions with the mean-squared displacement (MSD) given by $\langle{\tilde{\textbf{r}}}^{2}{\rangle}_{\rm st}=d/\beta$. This form is expected to hold at low enough activity. However, at large activity, even for non-chiral ABPs, the distribution deviates significantly from Gaussian behavior. These non-Gaussian departures can be characterized by exact calculations of fourth moment and kurtosis of displacement [67]. We compare the numerically obtained probability distributions with the Gaussian form supplemented by the exact MSD estimate shown below, to highlight the non-Gaussian departures (see solid lines in Figs. 2(d,e,f)).
In 2D, the distributions of non-chiral ABPs exhibited a circularly symmetric ring around the trap minimum, with particles climbing the trap potential until reaching a radial extent determined by the force balance condition: $\tilde{\rm{r}}^{0}_{\rm ac}={\rm Pe}/\beta$ [71, 67]. An equivalent estimate for chiral ABPs can be derived by analyzing the deterministic part of Langevin equations: $\dot{\tilde{x}}={\rm Pe}\cos\phi-\beta\tilde{x}$, $\dot{\tilde{y}}={\rm Pe}\sin\phi-\beta\tilde{y}$ and $\dot{\phi}=\Omega$. Solving these equations and obtaining the time-asymptotic expressions for $\tilde{x}(t)$ and $\tilde{y}(t)$, we find $\tilde{\rm{r}}_{\rm ac}={\rm Pe}/\sqrt{\beta^{2}+\Omega^{2}}$ (see Appendix-B). As we show later, this estimate agrees well with the off-center maxima of the probability distributions at high activity.
By applying the Fokker-Planck equation method (detailed derivation presented in Appendix-A), the first result we obtain is the MSD at steady state,
@@
The first term, denotes the equilibrium contribution in 2D, and the second term is entirely due to activity which increases as ${\rm Pe}^{2}$, but decreases quadratically with chirality as $\Omega^{-2}$. This result is consistent with observations in chiral active solid phases [24, 27, 72].
@@
II.2 Steady-state excess kurtosis and phase diagram
In the steady state, the deviation from Gaussian-like behavior is quantified by the excess kurtosis, defined as $\tilde{\cal{K}}_{\rm st}={\langle{{\tilde{\textbf{r}}}^{4}}\rangle_{\rm st}}/{\tilde{\mu}_{4}^{\rm st}}-1$, with $\tilde{\mu}_{4}=\langle{\tilde{\textbf{r}}}^{2}\rangle^{2}+2\langle\tilde{r}_{i}\tilde{r}_{j}\rangle^{2}$ giving the estimate of the fourth moment in a Gaussian process. Although vanishing $\tilde{\cal{K}}_{\rm st}$ may in principle arise from non-Gaussian distributions via specific cancellations, our numerical simulations confirm that the $\tilde{\cal{K}}_{\rm st}\approx 0$ regimes identified here are indeed Gaussian-like. In 2D, $\tilde{\mu}_{4}^{\rm st}=2\langle{\tilde{\textbf{r}}}^{2}\rangle_{\rm st}^{2}$ gives the following exact expression (see Appendix-A)
@@ II.2 Steady-state excess kurtosis and phase diagram
with $\tilde{\cal{K}}_{\rm st}$ vanishing linearly with $\beta$. In the small $\beta$ limit, $\tilde{\cal{K}}_{\rm st}$ behaves differently depending on $\Omega$: it is positive when $\Omega^{2}>7/5$ and negative when $\Omega^{2}<7/5$. As illustrated in Fig. 2(a), for instance, at $\Omega=5$ (blue square symbols) where $\Omega^{2}>7/5$, $\tilde{\cal{K}}_{\rm st}$ is initially positive at small $\beta$, but then becomes negative, following a $\beta^{2}$ dependence as the trap stiffness increases. After reaching a minimum at intermediate values of $\beta$, it begins to rise again, ultimately vanishing as $\sim-\beta^{-2}$ in the $\beta\to\infty$ limit.
@@
Phase diagram: Fig. 2(c) shows a phase diagram in the $\Omega-\beta$ plane, highlighting the transition between Gaussian (equilibrium-like) and non-Gaussian phases based on deviations from Gaussian statistics. While there is no true phase transition in this single-particle system, the diagram reveals a re-entrant crossover, shifting from Gaussian-like to non-Gaussian and back to Gaussian-like phases as trap stiffness increases. We identify three phases based on steady-state excess kurtosis: negative kurtosis defines the active phase (red) with a symmetric bimodal distribution with off-center peaks; vanishing kurtosis corresponds to the phase (blue) with a Gaussian-like distribution; and small positive kurtosis indicates another active phase (green) with a weakly heavy-tailed distribution, unique to chiral ABPs and absent in trapped achiral ABPs [67]. The solid line represents $\tilde{\cal{K}}_{\rm st}=0$, separating negative and positive kurtosis.
@@
Probability distributions: In Figs. 2(d-f), we present radial position distributions for the parameter values marked by Roman and Indo-Arabic numerals in the phase diagram of Fig. 2(c). The solid lines correspond to the Gaussian distribution $\rm p({\tilde{\textbf{r}}})=\exp{\left(-{{\tilde{\textbf{r}}}^{2}}/{\langle{\tilde{\textbf{r}}}^{2}\rangle_{\rm st}}\right)}/{\pi\langle{\tilde{\textbf{r}}}^{2}\rangle_{\rm st}}$ with $\langle{\tilde{\textbf{r}}}^{2}\rangle_{\rm st}$ for the corresponding parameter values. These help identify deviations from equilibrium-like Gaussian behavior.
@@
In Fig. 2(d), the position distribution is shown for varying trap stiffness $\beta=0.08~\mathrm{(i)},0.33~\mathrm{(ii)},10~\mathrm{(iii)},2\times 10^{4}~\mathrm{(iv)}$ with constant $\Omega=5$. The transition is observed from a weakly heavy-tailed active distribution ($\tilde{\cal K}_{\rm st}>0$) to a Gaussian-like, then to an active state with off-center peak ($\tilde{\cal K}_{\rm st}<0$), and finally back to a Gaussian-like state. Notably, for $\mathrm{(i)}~\beta=0.08$, although the excess kurtosis is positive, it is small ($\tilde{\cal{K}}_{\rm st}\approx 0.0084$), and the distribution appears nearly Gaussian. For $\mathrm{(ii)}~\beta=0.33$ on the zero excess kurtosis line, a Gaussian distribution is expected and observed. At $\mathrm{(iii)}~\beta=10$, a peak in the distribution away from the trap center, at $\tilde{\rm{r}}_{\rm ac}={\rm Pe}/\sqrt{\beta^{2}+\Omega^{2}}$, indicates an active state. At $\mathrm{(iv)}~\beta=2\times 10^{4}$, the distribution reverts to Gaussian.
@@
In Fig. 2(e), radial distributions across different chirality values ($\Omega$) with $\beta=10$ are shown, revealing transitions from active (off-center peak at $\tilde{\rm{r}}_{\rm ac}$) to Gaussian-like states. Fig. 2(f) at $\beta=0.15$ shows radial distributions at various $\Omega$ corresponding to small positive or negative kurtosis values with small departures from Gaussian.
@@
Analysis of the phase diagram: The phase diagram reveals an intermediate regime of $\beta$ supporting active dynamics marked by negative kurtosis. For $\beta>1$, persistence dominates over trap relaxation time allowing for active bimodal phase. Neglecting diffusion, the maximal radial reach of a trapped particle is $\tilde{\rm{r}}_{\rm ac}={\rm{Pe}}/\sqrt{\beta^{2}+\Omega^{2}}$ (see Appendix-B), while translational noise broadens the distribution over $\ell_{\rm eq}=\sqrt{2/\beta}$. This is the dimension-less form of $\ell_{\rm eq}\ell=\sqrt{2D/\mu}$. The off-center peak becomes negligible when $\tilde{\rm{r}}_{\rm ac}\lesssim\ell_{\rm eq}$, solution of the quadratic equation at $\tilde{\rm{r}}_{\rm ac}=\ell_{\rm eq}$ defines two bounds of the active bimodal phase,